OpenStudy (anonymous):

Differentiate WITHOUT using the quotient and product rule:

OpenStudy (anonymous):

\[y=\frac{ \sqrt{x}-2 }{ \sqrt[3]{x}}\]

OpenStudy (agent0smith):

Break it up into separate fractions first.

OpenStudy (anonymous):

yeah, i just did that..

OpenStudy (anonymous):

\[y=\frac{ \sqrt{x}}{ \sqrt[3]{x} }-\frac{ 2 }{ \sqrt[3]{x} }\] then?

OpenStudy (agent0smith):

Turn the radical signs into exponents (fractional exponents)

OpenStudy (agent0smith):

\[\Large y=\frac{x^\frac{ 1 }{ 2 }}{ x ^\frac{ 1 }{ 3 } }-\frac{ 2 }{ x ^\frac{ 1 }{ 3 } } = \] Now use exponent laws to simplify

OpenStudy (anonymous):